NEXT
User name J. Singh
Log entry time 05:39:02 on August21,2003
Entry number 112359
Followups:
keyword=owl shift summary, part 1
owl shift run summary
owl shift target summary
Beam was beautiful (again!) for the first half of the shift. At about 0400, there were problems with the helium pressure in one of the RF zones. A couple of times tonite we got really high deadtimes in the left arm. We "fixed" it, by cancelling, resetting, and downloading the left arm DAQ. There was no kcoda involved. In addition, we had another right arm DAQ "auto start failed." Right now, we are just waiting.
2357:0002 kin 10.17, ref cell = N2, 140psig, / IN
0002:0011 moving to Pol 3He
0011:0027 kin 10.17, Pol 3He, 0 deg / IN
0027:0029 BHWP = OUT
0029:0043 kin 10.17, Pol 3He, 0 deg / OUT
0043:0057 rotating to 270
0057:0112 kin 10.17, Pol 3He, 270 deg / OUT
0112:0141 changing to kin 10.18, rotate to 0 deg, move to ref cell, NMR = 38.6%
0141:0148 kin 10.18, ref cell = empty / OUT
0148:0152 filling w/ 142 psig N2, bleedthrough = 0.25%
0152:0158 kin 10.18, ref cell = N2 / OUT
0158:0209 moving to Pol 3He
0209:0227 kin 10.18, Pol 3He, 0 deg / OUT
0227:0230 BHWP = IN
0230:0246 kin 10.18, Pol 3He, 0 deg / IN
0246:0302 rotate to 270 deg
0302:0317 kin 10.18, Pol 3He, 270 deg / IN
0317:0344 changing to kin 10.19, rotate to 0 deg, move to ref cell
0344:0349 kin 10.19, ref cell = empty / IN
0349:0353 filling w/ 141 psig N2
0353:0358 kin 10.19, ref cell = N2 / IN
0358:0410 moving Pol 3He, bleedthrough = 0.73%
0410:0537 MCC called, problems, 112354, 112355, & 112356, moved to single foil to move Pol 3He cell out of the beam path.